Question Five Pink Ltd. sells a Product Exe with sales occurring evenly throughout the year. The annual demand for Product Exe is 30,000 units and an order for new inventory is placed each month. Each order costs KES 4,000 to place. The cost of holding Product Exe in inventory is KES 15 per unit per year. Buffer (safety) inventory equal to 40% of one month’s sales is maintained.

Required: Calculate the following values for Product Exe:

i. The total cost of the current ordering policy.

ii. The total cost of an ordering policy using the economic order quantity.

iii. The net cost or saving of introducing an ordering policy using the economic order quantity.
Answer #1

A.      
Number of orders per year =    12  
Ordering cost = Number of orders*order placing cost      
12*4000      
48000      
      
Safety stock = 40% of Monthly sales      
      
So Safety stock = 30000/12*40%=   1000  
Order Quantity = Annual demand/Number of orde      
30000/12=   2500  
Average Inventory = Safety stock +(order quantity/2)      
1000+(2500/2)      
2250      
Annual holding costs = holding cost per unit * Average Inventory      
15*2250      
33750      
      
Total cost = Annual order cost + annual holding cost      
48000+33750=   81750  
      
So Total cost of Current ordering policy is    81750  
      
B      
Annual usage (A)= 30000 units      
Ordering cost (o)= kes 4000      
Annual carrying cost (C)= kes 15      
Economic order quantity (EOQ) = √(2*A*O/C)      
√((2*30000*4000)/15)      
4000      
So EOQ quantity is 4000 units.      
      
Number of orders per year = 30000/4000=   8  
Ordering cost = Number of orders*order placing cost      
8*4000      
32000      
      
Safety stock = 40% of Monthly sales      
      
So Safety stock = 30000/12*40%=   1000  
      
      
Average Inventory = Safety stock +(order quantity/2)      
1000+(4000/2)      
3000      
Annual holding costs = holding cost per unit * Average Inventory      
15*3000      
45000      
      
Total cost = Annual order cost + annual holding cost      
32000+45000=   77000  
      
So Total cost of EOQ ordering policy is   77000  
      
c      
EOQ Quantity has less cost.       
Saving = cost of Current policy - Cost of EOQ policy      
81750-77000      
4750      
So net Savings due to choosing EOQ is 4750 kes      
      
please thumbs up